two cards are drawn from a standard deck of 52 cards without replacement. find the probability that the first card is an ace and the second is a 2.
Answer by jim_thompson5910(35256)

P(Ace then 2) = P(Ace)*P(2)

P(Ace then 2) = (4/52)*(4/51)

P(Ace then 2) = 16/2652

P(Ace then 2) = 4/663